OpenStudy (anonymous):

WILL MEDAL HELP PLEASE The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds Numerical Answers Expected!

OpenStudy (anonymous):

@pandaluvs

OpenStudy (anonymous):

*arrives in pink cape* i am here to saved the day!

OpenStudy (anonymous):

no, NO NO NO NO

OpenStudy (anonymous):

you betrayed me @A6020072944

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

he/she ASKED ME FOR HELP FIRST!!!

OpenStudy (anonymous):

ladies there plenty of room of my heart for the both of you XD

OpenStudy (anonymous):

and @AsoundofHeaven was helping me on my question so i could get to you!!

OpenStudy (anonymous):

THIS IS WARRRRRRR

OpenStudy (anonymous):

SORRY you can help me @pandaluvus

OpenStudy (anonymous):

̿’ ̿’\̵͇̿̿\з=(-_-)=ε/̵͇̿̿/’̿’̿ oh great time for me to hide behind the tree -hides behind tree-

OpenStudy (anonymous):

@AsoundofHeaven get back to my question.

OpenStudy (anonymous):

while i help @A6020072944

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

the equation given is 700A+1200R=1,000,000 why the cost should be fixed at 1,000,000 is beyond me, but anyway that is what it says then it says A=1000 so now the equation is 700×1000+1200R=1,000,000 or 700,000+1200R=1,000,000 which i would solve in two steps 1) subtract 7000,000 to get 1200R=300,000 2) then divide by 1200 to get R=250

OpenStudy (anonymous):

r is 250